what is the surface area of a cube with a side length of 2.1 ft

Answer by mananth(12269)

one side of cube = 2.1 ft.
alll sides are equal in a cube
area of one face of cube = 2.1*2.1 = 4.41ft^2
There are 6 faces for a cube
6*4.41 = 26.46 ft^2
surface area = 26.46 ft^2